What is the resonance frequency of a proton in a magnetic field of 14.6 T?
Ans = ______ MHz
Answer #1

gyromagnetic  ration for proton =42.58(MHz/T)

magnetic field field B0= 14.6 T

resonance frequency = * B0/

=42.58(MHz/T)*14.6 T/(2*3.14)

=621.668/(2*3.14))MHz

=98.99 MHZ